### 08 Jan 2024 v1

@@ 1 Introduction
The research in deep learning on AGI has witnessed explosive growth, particularly with the advent of LLMs like GPT4 OpenAI (2023), LLaMA Touvron et al. (2023b), accelerating the arrival of the AGI era. LLMs have achieved remarkable results in many downstream natural language tasks. Simultaneously, the development of multimodal large models, such as LLaVA Liu et al. (2023b), has sprung up like mushrooms after rain. There have also been outstanding achievements in the fields of vision, audio, 3D, and agent Lin et al. (2023a); Gong et al. (2023); Hong et al. (2023); Szot et al. (2023).
@@
However, despite the astonishing progress in the LLM of AGI, the outputs of the model still do not fully align with human expectations, which called hallucination Zhang et al. (2023). For instance, LLMs occasionally produce factually incorrect answers, LVLMs(Large Vision-Language Models) can sometimes generate responses that object error or something of their own fantasy, video language models also suffer actions inconsistency, and 3D or agent may encounter issues with hallucination across multiple perspectives. This problems hinders AGI’s path towards greater intelligence.
@@
Evaluating the hallucination error in AGI is also important area. By assessing LLMs, we gain deeper insights into their capabilities and limitations Huang et al. (2023); Chiang and yi Lee (2023); Chan et al. (2023a). Enhanced evaluation methodologies can significantly improve the interaction between humans and LLMs. This, in turn, has the potential to drive innovative designs and implementations in future interactions. Given the wide-ranging applications of LLMs, it is of utmost importance to ensure their safety and reliability, especially in sectors where safety is paramount, such as financial Sarmah et al. (2023) institutions and healthcare facilities Cai et al. (2023). There also has been a lot of progress in evaluating hallucinations with multimodal large models, including rule-based evaluation Li et al. (2023d); Lovenia et al. (2023), large model-based evaluation Wang et al. (2023a); Chan et al. (2023b), and human-based evaluation Xu et al. (2023); Liu and Wan (2023).
@@ 3.1 Training Data Distribution
The significance of training data on the efficacy of models is paramount. Both the quantity and quality of the data directly influence model’s distribution. The generalization ability of the model such as overfitting and underfitting to specific data distributions can adversely impact model function, occasionally leading to hallucinations in the model’s outputs. In the machine learning, Mougan et al. (2023)’s research highlighted that shifts in training data distribution markedly alter explanation characteristics. This findings suggest that data distribution changes can profoundly influence a model’s ultimate performance. Dong et al. (2023)’s work revealed that the quantity of compositional data significantly impacts model performance, whereas the ratio of composition bears minimal effect. Through integrating specialized domain data (mathematics and coding) into a general domain, the impact of varied data distributions on model performance was examined. The conclusion was that model capabilities are enhanced in scenarios with limited resources but diminish in high-resource contexts when compared to the capabilities derived from individual data sources. Furthermore, it was observed that as the model’s size increases, there is a corresponding amplification in performance gains in low-resource settings, particularly for mathematical and general competencies. Dziri et al. (2022)’s findings indicate that the proportion of hallucinations in generated responses is substantially higher compared to that in the training data. LIMA Zhou et al. (2023a) posit that the vast majority of knowledge in large language models is acquired during the pretraining phase, and that only a minimal amount of instruction tuning data is requisite for training these models to generate outputs of high quality. Li et al. (2020) stated that both overfitting and underfitting can impact the model’s performance.
@@ 4 Mitigation for AGI Hallucination
Leveraging a knowledge-base plays a crucial role in mitigating hallucinations during model inference and post-processing. ChatLaw Cui et al. (2023) proposes a strategy to counteract hallucination by augmenting the training process of the model and integrating four distinct modules during inference, namely ’consult’, ’reference’, ’self-suggestion’, and ’response’. Feng et al. (2023a) endeavor to instruct LLMs in querying relevant domain-specific knowledge from external knowledge graphs for answering specialized queries. Melz (2023) employs RAG (Retrieval Augmented Generation), Lewis et al. (2020), to enhance problem-solving efficacy. The Verify-then-Edit methodology Zhao et al. (2021) is focused on improving the factuality of predictions through post-editing reasoning chains, using external information sourced from Wikipedia. Peng et al. (2021) adopts the information retrieval paradigm, automating the identification of related entities in text-image pairs to boost model performance. Zhao et al. (2021) introduces an innovative approach that constructs a MMKG (multi-modal knowledge graph), linking visual objects with named entities and simultaneously capturing the relationships among these entities, aided by external knowledge obtained from the web. Jin et al. (2023) utilizes ClipBERT for extracting video-question features and derives frame-wise object-level external knowledge from a commonsense database to augment the performance of the model. Shapkin et al. (2023) introduces an end-to-end trainable architecture that incorporates a scalable entity retriever directly into the LLM decoder. This method significantly enhances the performance of code generation.
@@ 5 Evaluation for AGI Hallucination
The evaluation of AGI hallucinations is critically important, with substantial progress being made in AGI assessment in current research. Human evaluation of model-generated hallucinations is a time-consuming process, also humans themselves are susceptible to hallucinations. In the field of LLMs, Yu et al. (2023a); Sun et al. (2023a) have proposed the use of rule-based methods to detect model hallucinations, while Zha et al. (2023); Min et al. (2023); Liu and Wan (2023); Dhuliawala et al. (2023) have introduced automated assessment methods for detecting factual hallucinations, both of which are commendable contributions in automatically evaluate large model hallucinations. The occurrence of hallucinations in multimodal AGI is even more complex, encompassing not only misconceptions arising from language comprehension errors but also hallucinations related to the recognition of information in other modalities. In the context of LVLMs, POPE Li et al. (2023d) first proposed the use of discriminative methods to detect object hallucinations in images. Additionally, the work of Zhang et al. (2019); Yuan et al. (2021); Bai et al. (2023) highlights that methods based on LLMs can more effectively detect hallucinations. External knowledge bases also play a vital role in hallucination detection. In this section, we first introduce the benchmark for hallucination evaluation, then we categorize hallucination assessment into three types: rule-based, large model-based, and human-based.
@@
Benchmark   In both the traditional deep learning era and the era of large models, benchmarks play a crucial role in evaluating model illusions. Benchmarks in different fields, such as ChainPoll Friel and Sanyal (2023) and Bang et al. (2023), focus on assessing specific domain weaknesses, while benchmarks across different modalities evaluate various model skills. KoLA Yu et al. (2023a) introduced a knowledge-driven benchmark comprising 19 tasks, utilizing Wikipedia data along with continuously collected emerging corpora to construct more granular evaluations, addressing unseen data and evolving knowledge. UHGEval Liang et al. (2023) is an Unconstrained Hallucination Generation Evaluation benchmark for evaluating prominent Chinese language models. INVITE Ramakrishna et al. (2023) automatically generates invalid questions to assess large model illusions. Meanwhile, Head-to-Tail Sun et al. (2023a) uses a template-based method to automatically generate 18K QA (question-answer) pairs for model illusion assessment. DELUCIONQA Sadat et al. (2023) employs a QA system to automatically generate domain-specific questions and answers for detecting model illusions. Numerous benchmarks have also emerged in the fields of image and video, particularly for LVLMs such as MMBench Liu et al. (2023d), MM-Vet Yu et al. (2023b), SEED-Bench Li et al. (2023a), LVLM-eHub Xu et al. (2023), and AutoEval-Video Chen et al. (2023). MERLIM Villa et al. (2023) is an extensive database with over 279,000 image-question pairs, primarily focusing on identifying and analyzing cross-modal ’hallucination’ events in IT-LVLMs (Image-Text Language-Vision Language Models). In multimodal contexts, finer-grained captions can detect more nuanced hallucinations in image-text interactions, Gunjal et al. (2023) with a dataset of 16k finely annotated VQA examples. HALLUSIONBENCH Liu et al. (2023a) is the first to consider visual illusion and knowledge hallucination of LVLMs. RSGPT Hu et al. (2023) introduces a benchmark concerning Remote Sensing. In the video domain, benchmarks like AutoEval-Video Chen et al. (2023) , EgoSchema Mangalam et al. (2023), EvalCrafter Liu et al. (2023c), and MVBench Li et al. (2023c) have emerged. AutoEval-Video and MVBench establish video QA datasets, assessing Multi-modal Video from multiple perspectives. EgoSchema focuses on Very Long-form Video Language Understanding, while EvalCrafter and LLM4VG Feng et al. (2023b) emphasizes the assessment of Video Generation. Himakunthala et al. (2023) evaluates complex video reasoning tasks, and Liu and Wan (2023) evaluates the factuality datasets for video captioning.
@@
Furthermore, KoLA Yu et al. (2023a) initiative proposed the innovative Self-contrast Metric, rooted in ROUGE-L. This metric computes an average contrast score by juxtaposing results derived from knowledge-based prompts, standard prompts, and human responses, thereby effectively evaluating hallucinations in LLM-generated content.
For LVLMs, Shukor et al. (2023) adopted a unique approach, utilizing VQA accuracy, CIDEr Vedantam et al. (2014), and CHAIRs Rohrbach et al. (2018), specifically for appraising hallucinations within the COCO dataset. Object Hallucination, in particular, represents a formidable challenge in LVLMs. To address this problem, POPE Li et al. (2023d) , based on the co-occurrence principle of objects in images. POPE, a polling-based query methodology, is designed to scrutinize model hallucinations on objects through binary "is or not" questions. In a parallel development, NOPE Lovenia et al. (2023) leveraging LLMs to generate data featuring NegP (Negative Pronouns). NOPE’s approach is instrumental in evaluating object hallucination in these models.
@@ 6 Discourse for AGI Hallucination
Recent work has also shown that hallucinations are not entirely erroneous. Hallucinations play a role as adversarial examples to increase the robustness and creation of models, which need to produce hallucinations in contextually reasonable situations - akin to ’white lie’ in human life. Yao et al. (2023)’s research utilized weak semantic prompts and OOD prompts to elicit hallucinatory responses from LLMs. This approach led to a reassessment of hallucinations as a different perspective on adversarial examples. Zhang (2023) focused on the degree of faithfulness to reference knowledge in generated responses, striving for a balance between creativity and hallucination. Zhao et al. (2023) employed a method called Hallucination-Aware Direct Preference Optimization, using hallucinatory samples to optimize models through reinforcement learning, demonstrating this method’s effectiveness in mitigating hallucinations in LVLMs. Qiu et al. (2022) introduced an approach where a teacher iteratively generates synthetic training data based on the learner’s status, a process termed data hallucination teaching. Wu et al. (2023) developed a ’Hallucinator’ to generate additional positive image samples for enhanced contrast training. Fei et al. (2023) conceived a visual scene hallucination mechanism that dynamically creates pseudo visual scene graphs from textual scene graphs, significantly improving inference-time image-free unsupervised multimodal machine translation. Kulal et al. (2023) proposed inserting characters into scenes, enabling models to generate videos with both character and scene hallucinations, achieving harmonious composition and creativity. McKee et al. (2021) initially added video simulation augmentations to create hallucinated video data and then trained a tracker jointly on this hallucinated data and mined hard video examples. HALLUAUDIO Yu et al. (2023c) leverages a special audio format by hallucinating high-frequency and low-frequency parts as structured concepts for few-shot audio classification. Finally, Shah et al. (2023) pioneered the use of hallucinated latent positives in a skeleton-based CL (Contrastive Learning) framework. They also employed a MoCo-based framework that mixes latent space features of positives and negatives, better utilizing hallucinations in action generation tasks.
7 Talk about Future

### 17 Jan 2024 v2

@@ 1 Introduction
The research in deep learning on AGI has witnessed explosive growth, particularly with the advent of LLMs like GPT4 OpenAI (2023), LLaMA Touvron et al. (2023b), accelerating the arrival of the AGI era. LLMs have achieved remarkable results in many downstream natural language tasks. Simultaneously, the development of multimodal large models, such as LLaVA Liu et al. (2023b), has sprung up like mushrooms after rain. There have also been outstanding achievements in the fields of vision, audio, 3D, and agent Lin et al. (2023a); Gong et al. (2023); Hong et al. (2023); Szot et al. (2023a).
@@
However, despite the astonishing progress in the LLM of AGI, the outputs of the model still do not fully align with human expectations, which called hallucination Zhang et al. (2023c). For instance, LLMs occasionally produce factually incorrect answers, LVLMs(Large Vision-Language Models) can sometimes generate responses that object error or something of their own fantasy, video language models also suffer actions inconsistency, and 3D or agent may encounter issues with hallucination across multiple perspectives. This problems hinders AGI’s path towards greater intelligence.
@@
Evaluating the hallucination error in AGI is also important area. By assessing LLMs, we gain deeper insights into their capabilities and limitations Huang et al. (2023); Chiang and yi Lee (2023); Chan et al. (2023a). Enhanced evaluation methodologies can significantly improve the interaction between humans and LLMs. This, in turn, has the potential to drive innovative designs and implementations in future interactions. Given the wide-ranging applications of LLMs, it is of utmost importance to ensure their safety and reliability, especially in sectors where safety is paramount, such as financial Sarmah et al. (2023) institutions and healthcare facilities Cai et al. (2023). There also has been a lot of progress in evaluating hallucinations with multimodal large models, including rule-based evaluation Li et al. (2023f); Lovenia et al. (2023), large model-based evaluation Wang et al. (2023a); Chan et al. (2023b), and human-based evaluation Xu et al. (2023); Liu and Wan (2023).
@@ 3.1 Training Data Distribution
The significance of training data on the efficacy of models is paramount. Both the quantity and quality of the data directly influence model’s distribution Hammoudeh and Lowd (2022). The generalization ability of the model such as overfitting and underfitting to specific data distributions can adversely impact model function, occasionally leading to hallucinations in the model’s outputs. In the machine learning, Mougan et al. (2023)’s research highlighted that shifts in training data distribution markedly alter explanation characteristics. This findings suggest that data distribution changes can profoundly influence a model’s ultimate performance. Dong et al. (2023)’s work revealed that the quantity of compositional data significantly impacts model performance, whereas the ratio of composition bears minimal effect. Through integrating specialized domain data (mathematics and coding) into a general domain, the impact of varied data distributions on model performance was examined. The conclusion was that model capabilities are enhanced in scenarios with limited resources but diminish in high-resource contexts when compared to the capabilities derived from individual data sources. Furthermore, it was observed that as the model’s size increases, there is a corresponding amplification in performance gains in low-resource settings, particularly for mathematical and general competencies. Dziri et al. (2022)’s findings indicate that the proportion of hallucinations in generated responses is substantially higher compared to that in the training data. LIMA Zhou et al. (2023a) posit that the vast majority of knowledge in large language models is acquired during the pretraining phase, and that only a minimal amount of instruction tuning data is requisite for training these models to generate outputs of high quality. Li et al. (2020) stated that both overfitting and underfitting can impact the model’s performance.
@@ 4 Mitigation for AGI Hallucination
Area Stage Works
Language Data
Brown et al. (2020), Ouyang et al. (2022), Touvron et al. (2023c), Penedo et al. (2023)
Train & SFT & RLHF
Touvron et al. (2023a), OpenAI (2023), Rafailov et al. (2023)
Inference & Post-hoc
Krishna et al. (2023), Cui et al. (2023), Feng et al. (2023a), Melz (2023), Zhao et al. (2021), Shapkin et al. (2023)
Visual Language Data
Wang et al. (2023c)
Train & SFT & RLHF
Liu et al. (2023b), Yoon et al. (2022), Sun et al. (2023b)
Inference & Post-hoc
Agarwal et al. (2021), Peng et al. (2021), Zhao et al. (2021), Zhou et al. (2023b)
Video Language Data
Li et al. (2023b)
Train & SFT & RLHF
Li et al. (2023c), Jin et al. (2023a), Yu et al. (2023d), Zhang et al. (2023b), Shukor et al. (2023a), Jin et al. (2023b)
Inference & Post-hoc
Jin et al. (2023c),Yin et al. (2023)
Audio Language Train & SFT & RLHF
Kanda et al. (2023), Serai et al. (2022), Sridhar et al. (2023), Doh et al. (2023), Gong et al. (2023)
Inference & Post-hoc
Lyu et al. (2023)
3D & Agent Train & SFT & RLHF
Mu et al. (2022), Li et al. (2023e), Gong et al. (2022), Ren et al. (2023), Szot et al. (2023b), Xia et al. (2023)
Table 2: Overview of Migitation for AGI Hallucination.
@@
Leveraging a knowledge-base plays a crucial role in mitigating hallucinations during model inference and post-processing. ChatLaw Cui et al. (2023) proposes a strategy to counteract hallucination by augmenting the training process of the model and integrating four distinct modules during inference, namely ’consult’, ’reference’, ’self-suggestion’, and ’response’. Feng et al. (2023a) endeavor to instruct LLMs in querying relevant domain-specific knowledge from external knowledge graphs for answering specialized queries. Melz (2023) employs RAG (Retrieval Augmented Generation), Lewis et al. (2020), to enhance problem-solving efficacy. The Verify-then-Edit methodology Zhao et al. (2021) is focused on improving the factuality of predictions through post-editing reasoning chains, using external information sourced from Wikipedia. Peng et al. (2021) adopts the information retrieval paradigm, automating the identification of related entities in text-image pairs to boost model performance. Zhao et al. (2021) introduces an innovative approach that constructs a MMKG (multi-modal knowledge graph), linking visual objects with named entities and simultaneously capturing the relationships among these entities, aided by external knowledge obtained from the web. Jin et al. (2023c) utilizes ClipBERT for extracting video-question features and derives frame-wise object-level external knowledge from a commonsense database to augment the performance of the model. Shapkin et al. (2023) introduces an end-to-end trainable architecture that incorporates a scalable entity retriever directly into the LLM decoder. This method significantly enhances the performance of code generation.
Area Eval-Mod Works
Language Benchmark
Friel and Sanyal (2023), Bang et al. (2023), Yu et al. (2023b), Liang et al. (2023), Ramakrishna et al. (2023),Sun et al. (2023a),Sadat et al. (2023)
Rule-Based
Lin (2004), Papineni et al. (2002), Sun et al. (2023a), Min et al. (2023), Yu et al. (2023b)
Large Model-Based
Zhang et al. (2019), Yuan et al. (2021), Zha et al. (2023), Chan et al. (2023b),Sarmah et al. (2023), Chan et al. (2023a)
Human-Based
Chiang and yi Lee (2023), Min et al. (2023)
Visual Language Benchmark
Bang et al. (2023), Liu et al. (2023d), Yu et al. (2023c), Li et al. (2023a), Xu et al. (2023),Villa et al. (2023), Gunjal et al. (2023), Liu et al. (2023a), Hu et al. (2023)
Rule-Based
Shukor et al. (2023b), Vedantam et al. (2014), Rohrbach et al. (2018), Li et al. (2023f), Lovenia et al. (2023)
Large Model-Based
Zhai et al. (2023a), Bai et al. (2023)
Human-Based
Guan et al. (2023), Xu et al. (2023), fei Yin et al. (2023)
Video Language Benchmark
Bang et al. (2023), Chen et al. (2023a), Chen et al. (2023a), Mangalam et al. (2023), Liu et al. (2023c), Li et al. (2023d), Feng et al. (2023b), Himakunthala et al. (2023), Liu and Wan (2023), Grauman et al. (2023)
Rule-Based
Vedantam et al. (2014), Rohrbach et al. (2018)
Large Model-Based
Zhang et al. (2023a), Feng et al. (2023b), Chen et al. (2023b)
Human-Based
Ma et al. (2023), Chen et al. (2023c)
Audio Language Benchmark
Behera et al. (2023), de Seyssel et al. (2023)
Rule-Based
Yu et al. (2023a)
3D & Agent Benchmark
Behera et al. (2023), Chen et al. (2023c)
Large Model-Based
Yang et al. (2023)
Human-Based
fei Yin et al. (2023)
Table 3: Overview of Evaluation for AGI Hallucination.
@@ 5 Evaluation for AGI Hallucination
The evaluation of AGI hallucinations is critically important, with substantial progress being made in AGI assessment in current research. Human evaluation of model-generated hallucinations is a time-consuming process, also humans themselves are susceptible to hallucinations. In the field of LLMs, Yu et al. (2023b); Sun et al. (2023a) have proposed the use of rule-based methods to detect model hallucinations, while Zha et al. (2023); Min et al. (2023); Liu and Wan (2023); Dhuliawala et al. (2023) have introduced automated assessment methods for detecting factual hallucinations, both of which are commendable contributions in automatically evaluate large model hallucinations. The occurrence of hallucinations in multimodal AGI is even more complex, encompassing not only misconceptions arising from language comprehension errors but also hallucinations related to the recognition of information in other modalities. In the context of LVLMs, POPE Li et al. (2023f) first proposed the use of discriminative methods to detect object hallucinations in images. Additionally, the work of Zhang et al. (2019); Yuan et al. (2021); Bai et al. (2023) highlights that methods based on LLMs can more effectively detect hallucinations. External knowledge bases also play a vital role in hallucination detection. In this section, we first introduce the benchmark for hallucination evaluation, then we categorize hallucination assessment into three types: rule-based, large model-based, and human-based.
@@
Benchmark   In both the traditional deep learning era and the era of large models, benchmarks play a crucial role in evaluating model illusions. Benchmarks in different fields, such as ChainPoll Friel and Sanyal (2023) and Bang et al. (2023), focus on assessing specific domain weaknesses, while benchmarks across different modalities evaluate various model skills. KoLA Yu et al. (2023b) introduced a knowledge-driven benchmark comprising 19 tasks, utilizing Wikipedia data along with continuously collected emerging corpora to construct more granular evaluations, addressing unseen data and evolving knowledge. UHGEval Liang et al. (2023) is an Unconstrained Hallucination Generation Evaluation benchmark for evaluating prominent Chinese language models. INVITE Ramakrishna et al. (2023) automatically generates invalid questions to assess large model illusions. Meanwhile, Head-to-Tail Sun et al. (2023a) uses a template-based method to automatically generate 18K QA (question-answer) pairs for model illusion assessment. DELUCIONQA Sadat et al. (2023) employs a QA system to automatically generate domain-specific questions and answers for detecting model illusions. Numerous benchmarks have also emerged in the fields of image and video, particularly for LVLMs such as MMBench Liu et al. (2023d), MM-Vet Yu et al. (2023c), SEED-Bench Li et al. (2023a), LVLM-eHub Xu et al. (2023), and AutoEval-Video Chen et al. (2023a). MERLIM Villa et al. (2023) is an extensive database with over 279,000 image-question pairs, primarily focusing on identifying and analyzing cross-modal ’hallucination’ events in IT-LVLMs (Image-Text Language-Vision Language Models). In multimodal contexts, finer-grained captions can detect more nuanced hallucinations in image-text interactions, Gunjal et al. (2023) with a dataset of 16k finely annotated VQA examples. HALLUSIONBENCH Liu et al. (2023a) is the first to consider visual illusion and knowledge hallucination of LVLMs. RSGPT Hu et al. (2023) introduces a benchmark concerning Remote Sensing. In the video domain, benchmarks like AutoEval-Video Chen et al. (2023a) , EgoSchema Mangalam et al. (2023), EvalCrafter Liu et al. (2023c), and MVBench Li et al. (2023d) have emerged. AutoEval-Video and MVBench establish video QA datasets, assessing Multi-modal Video from multiple perspectives. EgoSchema focuses on Very Long-form Video Language Understanding, while EvalCrafter and LLM4VG Feng et al. (2023b) emphasizes the assessment of Video Generation. Himakunthala et al. (2023) evaluates complex video reasoning tasks, and Liu and Wan (2023) evaluates the factuality datasets for video captioning. Grauman et al. (2023) present an exocentric video of skilled human activities as multimodal multiview video dataset and benchmark challenge. Behera et al. (2023) used LLM to generate expansive, high-quality Audio Question Answering datasets, contributing significantly to the progression of Audio research.
@@
Furthermore, KoLA Yu et al. (2023b) initiative proposed the innovative Self-contrast Metric, rooted in ROUGE-L. This metric computes an average contrast score by juxtaposing results derived from knowledge-based prompts, standard prompts, and human responses, thereby effectively evaluating hallucinations in LLM-generated content.
For LVLMs, Shukor et al. (2023b) adopted a unique approach, utilizing VQA accuracy, CIDEr Vedantam et al. (2014), and CHAIRs Rohrbach et al. (2018), specifically for appraising hallucinations within the COCO dataset. Object Hallucination, in particular, represents a formidable challenge in LVLMs. To address this problem, POPE Li et al. (2023f) , based on the co-occurrence principle of objects in images. POPE, a polling-based query methodology, is designed to scrutinize model hallucinations on objects through binary "is or not" questions. In a parallel development, NOPE Lovenia et al. (2023) leveraging LLMs to generate data featuring NegP (Negative Pronouns). NOPE’s approach is instrumental in evaluating object hallucination in these models.
@@ 6 Discourse for AGI Hallucination
Recent work has also shown that hallucinations are not entirely erroneous. Hallucinations play a role as adversarial examples to increase the robustness and creation of models, which need to produce hallucinations in contextually reasonable situations - akin to ’white lie’ in human life. Yao et al. (2023)’s research utilized weak semantic prompts and OOD prompts to elicit hallucinatory responses from LLMs. This approach led to a reassessment of hallucinations as a different perspective on adversarial examples. Zhang (2023) focused on the degree of faithfulness to reference knowledge in generated responses, striving for a balance between creativity and hallucination. Zhao et al. (2023) employed a method called Hallucination-Aware Direct Preference Optimization, using hallucinatory samples to optimize models through reinforcement learning, demonstrating this method’s effectiveness in mitigating hallucinations in LVLMs. Qiu et al. (2022) introduced an approach where a teacher iteratively generates synthetic training data based on the learner’s status, a process termed data hallucination teaching. Wu et al. (2023) developed a ’Hallucinator’ to generate additional positive image samples for enhanced contrast training. Fei et al. (2023) conceived a visual scene hallucination mechanism that dynamically creates pseudo visual scene graphs from textual scene graphs, significantly improving inference-time image-free unsupervised multimodal machine translation. Kulal et al. (2023) proposed inserting characters into scenes, enabling models to generate videos with both character and scene hallucinations, achieving harmonious composition and creativity. McKee et al. (2021) initially added video simulation augmentations to create hallucinated video data and then trained a tracker jointly on this hallucinated data and mined hard video examples. HALLUAUDIO Yu et al. (2023e) leverages a special audio format by hallucinating high-frequency and low-frequency parts as structured concepts for few-shot audio classification. Finally, Shah et al. (2023) pioneered the use of hallucinated latent positives in a skeleton-based CL (Contrastive Learning) framework. They also employed a MoCo-based framework that mixes latent space features of positives and negatives, better utilizing hallucinations in action generation tasks.
@@
7 Talking about Future